# Cluster-based Graph Collaborative Filtering

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recommendation systems have become essential for various online platforms. It not only assists users in navigating through overwhelming volumes of information but also serves as a powerful tool for service providers to increase customer retention and revenue. Within this field, Collaborative Filtering (CF)-based models have achieved significant advancements in learning the user and item representations by effectively exploiting the historical user-item interactions (Bell and Koren, 2007; He et al., 2017; Wu et al., 2016; Xue et al., 2017). For instance, Matrix Factorization (MF) decomposes the user-item interaction matrices into lower-dimensional latent vectors and predicts the user preference with the inner product (Bell and Koren, 2007). By utilizing nonlinear neural networks, NeuMF (He et al., 2017) can model the complex interaction behaviors, yielding better recommendation performance. More recently, Graph Convolution Networks (GCNs)-based models (van den Berg et al., 2018; Wang et al., 2019b, a; He et al., 2020; Liu et al., 2022; Wang et al., 2020; Liu et al., 2021; Fan et al., 2022; Wei et al., 2023) offer a cutting-edge approach in recommendation systems by leveraging high-order connectivities based on the topological structure of the interaction data. GCN-based models learn node representations rely on iteratively aggregating information from neighboring nodes. Many methods have demonstrated that exploring additional information based on graph topology structure can efficiently enhance user and item representation, ultimately improving recommendation performance. For example, NGCF (Wang et al., 2019b) incorporates high-order connectivity of user-item interactions in the user-item bipartite graph for representation learning in recommendation. With careful experimental studies, He et al. (He et al., 2020) simplified NGCF and achieved better performance by removing the feature transformation and the non-linear activation module. This is because these two components do not contribute positively to the final performance, and may even have a negative effect due to their impact on training difficulty.
@@
Despite the notable success of GCNs in recommendation systems, two key challenges persist in the information aggregation process of neighboring nodes. Firstly, noisy information from unreliable neighboring nodes harms the representation of the target nodes (Liu et al., 2021; Fan et al., 2022). Most existing GCN-based recommendation models are based on a user-item graph where interactions link users and items. They generally assume that embedding learning could take advantage of collaborative signals from high-order neighbors. Thus, they aggregate the messages passing from neighboring nodes to form the target user/item representation without distinguishing the high-order neighbors. As a result, the noisy information from unreliable users and items is also involved in the representation learning of a target node, resulting in sub-optimal performance. For instance, user $u_{1}$ has a strong preference for mystery novels. However, a high-order neighboring user $u_{2}$, who is loosely connected to $u_{1}$, prefers gardening books. In GCN-based models, this noisy information about gardening books could be delivered to the representation of $u_{1}$. Some efforts have been made to enhance the quality of information delivered from the neighbor nodes. For example, IMP-GCN (Liu et al., 2021) devises an interest-aware strategy to only pass messages from user nodes with common interest and their associated item nodes when performing high-order graph convolution. To capture the adaptive reliability of the interactions between users and items, GTN (Fan et al., 2022) introduces a graph trend collaborative filtering technique to achieve the desired adaptive smoothness property for the interaction graph. Secondly, GCN-based recommendation methods suffer from the over-smoothing problem (Wang et al., 2019b; Chen et al., 2020; He et al., 2020; Liu et al., 2021). Most GCN-based recommendation models achieve their peak performance by stacking only a few layers, typically 2 or 3. Increasing the number of layers leads to a significant drop in performance, as the graph convolution operation smooths the graph Laplacian and makes node representations indistinguishable after multiple convolutions (Liu et al., 2021). For example, Chen et al. (Chen et al., 2020) found that stacking more layers leads to over-smoothing, causing user and item representations to become more similar in recommendation. To alleviate this problem, they simplified the network structure by removing the non-linearities and introduced a residual network structure, which significantly improved the accuracy of recommendations. To maintain the uniqueness of the users by stacking more graph convolution layers, Liu et al. (Liu et al., 2021) grouped users based on their shared interests and performed high-order graph convolution on subgraphs.
@@
Although the aforementioned methods show significant advancements, they overlooked the users’ multiple interests in high-order graph convolution operations. As demonstrated in  (Liu et al., 2021), users receive more valuable information from other users with common interests and their associated items. However, users often exhibit diverse and overlapping interests across a range of topics or item categories (e.g., movies of different genres). For example, a user might prefer both science fiction and romantic movies. Similarly, a movie is probably categorized into both science fiction and action. As illustrated in Fig. 1, a user node could be classified into multiple clusters representing different interests. Specifically, user $u_{3}$ has a common interest with user $u_{1}$ in cluster 1 while also sharing an interest with user $u_{2}$ within cluster 2. Yet, there is no common interest between $u_{1}$ and $u_{2}$. Similarly, items can be classified into multiple clusters based on their characteristics. Therefore, it is necessary to jointly consider the multiple interests of each user and the common interest among users for capturing accurate and comprehensive representations of users and items, particularly when performing high-order graph convolution operations.
@@
Motivated by the above consideration, we propose a novel recommendation model termed Cluster-based Graph Collaborative Filtering (ClusterGCF). This model performs high-order graph convolution operations on cluster-specific graphs, which are constructed by capturing the multiple interests of users and identifying the common interests among them. To construct the cluster-specific graphs, we introduce an unsupervised and optimizable soft node clustering approach that classifies all the user and item nodes into multiple clusters by leveraging their embeddings and the graph topology information. Based on the soft node clustering results and the topology of the user-item interaction graph, we assign the nodes with probabilities for different clusters to construct the cluster-specific graphs. Since the embeddings of nodes are propagated within cluster-specific graphs during the graph convolution process, our model can filter out the negative information while capturing more valuable information from high-order neighboring nodes. Besides, the proposed ClusterGCF maintains the uniqueness of node embeddings while stacking more graph convolution layers because it can benefit from higher-order neighboring nodes in cluster-specific graphs. Our model’s effectiveness was validated through extensive experimentation on four publicly available datasets. Our ClusterGCF results demonstrate superior performance compared to the state-of-the-art GCN-based recommendation methods. This indicates that our model can benefit from higher-order neighbors by classifying nodes into multiple clusters. Furthermore, the results also verified the effectiveness of our proposed soft node clustering approach. We released the codes and involved parameter settings to facilitate others to repeat this work  [^1] https://github.com/zhao254014/ClusterGCF..
@@ 2.1. Collaborative Filtering
Collaborative Filtering (CF) techniques (Koren et al., 2009; Rendle et al., 2009; Wang et al., 2019a; He et al., 2017; Wang et al., 2019b; He et al., 2020; Wang et al., 2022) are widely used in modern recommendation systems and play an essential role in suggesting items to users. Their core idea is to learn users’ and items’ representations ($aka.$ embedding vectors) by exploiting the interaction behavior between them. Among them, Matrix Factorization (MF) (Koren et al., 2009) is a typical method that learns user and item embeddings by reconstructing the user-item interaction matrix. Due to its powerful capabilities, deep learning is widely used in recommendations to learn user and item embeddings (Cheng et al., 2016; Xue et al., 2017; Li et al., 2019). For example, Wide & Deep (Cheng et al., 2016) adopts deep neural networks to learn dense low-dimensional embedding for sparse features; DMF (Xue et al., 2017) learns a common low-dimensional space for user and item representations using a deep structure learning architecture that combines explicit ratings and implicit feedback. To improve the recommendation performance, the Deep learning technique is further employed to capture fine-grained user preferences. For example, NAIS (He et al., 2018) employs an attention network to estimate the importance of interacted items for user preference; MAML (Liu et al., 2019) introduces a multimodal attentive approach to capture user’s diverse preferences. After obtaining the embedding vectors for both the user and the item, a preferred score is calculated between them using the dot product (Rendle et al., 2009; Koren et al., 2009). Due to the intricate nature of user-item interactions, deep learning techniques are also employed as the interaction prediction function in recommendation systems (He et al., 2017).
@@ 2.2. GCN-based Recommendation Models
Early recommendation approaches used random walks on graphs to leverage the high-order proximity between users and items, inferring indirect user preference (Christoffel et al., 2015; Fouss et al., 2007). Some recent approaches leverage the bipartite graph structure of user-item interactions to enhance these relationships (Yu et al., 2018). Furthermore, they probe further to uncover collaborative relations, such as similarities between users and items (Chen et al., 2019). The efficacy of these methods is closely tied to the quality of interactions sampled through the random walks, as they rely heavily on these samples for model training. In recent years, Graph Convolutional Networks (GCNs) have garnered significant attention in the realm of recommendation systems. Their rise in popularity can be attributed to their adeptness at learning node representations from non-Euclidean structures (van den Berg et al., 2018; Wang et al., 2019b, a, 2020; Huang et al., 2021; He et al., 2020; Liu et al., 2021; Wang and Leskovec, 2021; Lin et al., 2022; Cheng et al., 2023; Wei et al., 2023; Liu et al., 2023). For example, NGCF (Wang et al., 2019b) propagates embeddings on the user-item interaction graph to harness the benefits of the high-order proximity. Drawing inspiration from studies focused on simplifying GCNs (Wu et al., 2019), researchers also critically scrutinize the intricacies of GCN-based recommendation models. Notably, He et al. (He et al., 2020) observed that the two commonly adopted design elements—feature transformation and nonlinear activation—did not contribute positively to the final performance. Consequently, they introduced LightGCN, a model that omits these two components and yields a marked improvement in recommendation performance. To achieve more efficient recommendations, the UltraGCN (Mao et al., 2021) model further simplifies graph convolutional networks by omitting embedding propagation layers. Additionally, it incorporates higher-order relationships through user-user graphs, leading to better performance and shorter running time compared to LightGCN. To further improve the neural graph collaborative filtering, NCL (Lin et al., 2022) explicitly incorporates the structural neighbors and semantic neighbors into contrastive pairs. Since the GCN-based recommendation methods suffer from the over-smoothing problem, IMP-GCN (Liu et al., 2021) refines user and item embeddings by conducting high-order graph convolutions within subgraphs. It can benefit from higher-order neighboring nodes by filtering out the noisy information from them. To learn more effective and robust representations, Fan et al. (Fan et al., 2022) proposed GTN, which can adaptively determine the reliability of interactions and effectively reduce unreliable interaction behaviors.
@@ 3.2. ClusterGCF Model
In this section, we present Cluster-based Graph Collaborative Filtering (ClusterGCF) model, along with its training methodology for recommendations. As illustrated in Figure 2, ClusterGCF performs the first-order graph convolution over the original user-item interaction graph  [^2] Note that the first-order graph convolution is adopted as implemented in IMP-GCN (Liu et al., 2021). and the high-order graph convolution on the cluster-specific graphs to form the final representation of users and items. To construct cluster-specific graphs, we propose a soft node clustering approach to group users and items into multiple clusters. Next, based on the results of clustering, we assign a probability score to each node in the original user-item interaction graph for each cluster. This allows us to generate a graph that is specific to each cluster and shows the connections between users and items within that cluster. In the following, we will provide details of the ClusterGCF.
@@ 3.2.2. High-Order Graph Convolution
Cluster-Specific Graphs Combination. The $k-$th layer representations of user $u$ and item $i$ ($k\geq 2$) are the combinations of their embeddings in multiple cluster-specific graphs:
@@
where $\bm{e_{u}^{(k)}}$ and $\bm{e_{i}^{(k)}}$ denote the features of user $u$ and item $i$ aggregated from the $k$-th layer neighbor nodes, respectively.
@@
After performing $k$ layers of graph convolution, the final embeddings of user $u$ and item $i$ are obtained by combing their embeddings across all layers of the ClusterGCF:
@@ 3.3. Model Training
In this work, our goal is to recommend a list of items to the target users by computing their preference score. For the optimization, we adopt the pairwise learning method, which is similar to the approach employed in previous works (Liu et al., 2021). To train our ClusterGCF model, we construct a triplet of $\{u,i^{+},i^{-}\}$, where $i^{+}$ represents a positive item and $i^{+}$ denotes a negative item. The objective function for ClusterGCF is defined as:

### v2

@@ Abstract.
ccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recommendation systems have become essential for various online platforms. It not only assists users in navigating through overwhelming volumes of information but also serves as a powerful tool for service providers to increase customer retention and revenue. Within this field, Collaborative Filtering (CF)-based models have achieved significant advancements in learning the user and item representations by effectively exploiting the historical user-item interactions (Bell and Koren, 2007; He et al., 2017; Wu et al., 2016; Xue et al., 2017). For instance, Matrix Factorization (MF) decomposes the user-item interaction matrices into lower-dimensional latent vectors and predicts the user preference with the inner product (Bell and Koren, 2007). By utilizing nonlinear neural networks, NeuMF (He et al., 2017) can model the complex interaction behaviors, yielding better recommendation performance. More recently, Graph Convolution Networks (GCNs)-based models (van den Berg et al., 2018; Wang et al., 2019b, a; He et al., 2020; Liu et al., 2022; Wang et al., 2020; Liu et al., 2021; Fan et al., 2022; Wei et al., 2023) offer a cutting-edge approach in recommendation systems by leveraging high-order connectivities based on the topological structure of the interaction data. These models obtain the final node representations ($a.k.a$ user and item embeddings) by iteratively aggregating information from neighboring nodes. Many methods have demonstrated that exploring additional information based on graph topology structure can efficiently enhance user and item representation, ultimately improving recommendation performance. For example, NGCF (Wang et al., 2019b) incorporates high-order connectivity of user-item interactions in the user-item bipartite graph for representation learning in recommendation. With careful experimental studies, He et al. (He et al., 2020) simplified NGCF and achieved better performance by removing the feature transformation and the non-linear activation module. This is because these two components do not contribute positively to the final performance, and may even have a negative effect due to their impact on training difficulty.
@@
Despite the notable success of GCNs in recommendation systems, two key challenges persist in the information aggregation process of neighboring nodes. Firstly, noisy information from unreliable neighboring nodes harms the representation of the target nodes (Liu et al., 2021; Fan et al., 2022). Most existing GCN-based recommendation models are based on a user-item graph where interactions link users and items. They generally assume that embedding learning could take advantage of collaborative signals from high-order neighbors. Thus, they aggregate the messages passing from neighboring nodes to form the target user/item representation without distinguishing the high-order neighbors. As a result, the noisy information from unreliable users and items is also involved in the representation learning of a target node, resulting in sub-optimal performance. For instance, user $u_{1}$ has a strong preference for mystery novels. However, a high-order neighboring user $u_{2}$, who is loosely connected to $u_{1}$, prefers gardening books. In GCN-based models, this noisy information about gardening books could be delivered to the representation of $u_{1}$. Some efforts have been made to enhance the quality of information delivered from the neighbor nodes. For example, IMP-GCN (Liu et al., 2021) devises an interest-aware strategy to only pass messages from user nodes with common interest and their associated item nodes when performing high-order graph convolution. To capture the adaptive reliability of the interactions between users and items, GTN (Fan et al., 2022) introduces a graph trend collaborative filtering technique to achieve the desired adaptive smoothness property for the interaction graph. Secondly, GCN-based recommendation methods suffer from the over-smoothing problem (Wang et al., 2019b; Chen et al., 2020; He et al., 2020; Liu et al., 2021). Most GCN-based recommendation models achieve their peak performance by stacking only a few layers, typically 2 or 3. Increasing the number of layers leads to a significant drop in performance, as the graph convolution operation smooths the graph Laplacian and makes node representations indistinguishable after multiple convolutions (Liu et al., 2021). For example, Chen et al. (Chen et al., 2020) found that stacking more layers leads to over-smoothing, causing user and item representations to become more similar in recommendation. To alleviate this problem, they simplified the network structure by removing the non-linearities and introduced a residual network structure, which significantly improved the accuracy of recommendations. To maintain the uniqueness of the users by stacking more graph convolution layers, Liu et al. (Liu et al., 2021) grouped users based on their shared interests and performed high-order graph convolution on subgraphs. Furthermore, as demonstrated in (Nembrini et al., 2022), detecting the users within each community who share similar tastes can improve recommendation accuracy.
@@
Given these challenges, previous works generally ignore high-order connections within the graph, overlooking users’ multiple interests in high-order graph convolution operations. Indeed, in  (Liu et al., 2021), the authors demonstrated that users receive more valuable information from other users with common interests and their associated items. As illustrated in Fig. 1(a) and (b), to identify the users with common interests, they classify users into different groups using a hard clustering method. However, users often exhibit diverse and overlapping interests across a range of topics or item categories (e.g., movies of different genres). For example, a user might prefer both science fiction and romantic movies. Similarly, a movie is probably categorized into both science fiction and action. The hard clustering fails to capture this overlap, resulting in limited utilization of information. In contrast, with soft clustering, a user could be classified into multiple clusters representing different interests. As shown in Fig. 1(c), user $u_{3}$ has a common interest with user $u_{1}$ in cluster 1 while also sharing an interest with user $u_{2}$ within cluster 2. Yet, there is no common interest between $u_{1}$ and $u_{2}$. Similarly, items can be classified into multiple clusters based on their characteristics. To fully exploit the valuable information from neighboring nodes, it is necessary to jointly consider the multiple interests of each user and the common interest among users for capturing accurate and comprehensive representations of users and items, particularly when performing high-order graph convolution operations.
@@
Motivated by the above consideration, we propose a novel recommendation model termed Cluster-based Graph Collaborative Filtering (ClusterGCF). This model performs high-order graph convolution operations on cluster-specific graphs, which are constructed by capturing the multiple interests of users and identifying the common interests among them. To construct the cluster-specific graphs, we introduce an unsupervised and optimizable soft node clustering approach that classifies all the user and item nodes into multiple clusters by leveraging their embeddings and the graph topology information. Based on the soft node clustering results and the topology of the user-item interaction graph, we assign the nodes with probabilities for different clusters to construct the cluster-specific graphs. Since the embeddings of nodes are propagated within cluster-specific graphs during the graph convolution process, our model can filter out the negative information while capturing more valuable information from high-order neighboring nodes. Besides, the proposed ClusterGCF maintains the uniqueness of node embeddings while stacking more graph convolution layers because it can benefit from higher-order neighboring nodes assigned with different probabilities in cluster-specific graphs. As a result, the over-smoothing problem could be alleviated. Our model’s effectiveness was validated through extensive experimentation on four publicly available datasets. Our ClusterGCF results demonstrate superior performance compared to the state-of-the-art GCN-based recommendation methods. This indicates that our model can benefit from higher-order neighbors by classifying nodes into multiple clusters. Furthermore, the results also verified the effectiveness of our proposed soft node clustering approach. We released the codes and involved parameter settings to facilitate others to repeat this work  [^1] https://github.com/zhao254014/ClusterGCF..
@@ 2.1. Collaborative Filtering
Collaborative Filtering (CF) techniques (Koren et al., 2009; Rendle et al., 2009; Wang et al., 2019a; He et al., 2017; Wang et al., 2019b; He et al., 2020; Wang et al., 2022; Liu et al., 2023b) are widely used in modern recommendation systems and play an essential role in suggesting items to users. Their core idea is to learn users’ and items’ representations ($aka.$ embedding vectors) by exploiting the interaction behavior between them. Among them, Matrix Factorization (MF) (Koren et al., 2009) is a typical method that learns user and item embeddings by reconstructing the user-item interaction matrix. Due to its powerful capabilities, deep learning is widely used in recommendations to learn user and item embeddings (Cheng et al., 2016; Xue et al., 2017; Li et al., 2019). For example, Wide & Deep (Cheng et al., 2016) adopts deep neural networks to learn dense low-dimensional embedding for sparse features; DMF (Xue et al., 2017) learns a common low-dimensional space for user and item representations using a deep structure learning architecture that combines explicit ratings and implicit feedback. To improve the recommendation performance, the Deep learning technique is further employed to capture fine-grained user preferences. For example, NAIS (He et al., 2018) employs an attention network to estimate the importance of interacted items for user preference; MAML (Liu et al., 2019) introduces a multimodal attentive approach to capture user’s diverse preferences. After obtaining the embedding vectors for both the user and the item, a preferred score is calculated between them using the dot product (Rendle et al., 2009; Koren et al., 2009). Due to the intricate nature of user-item interactions, deep learning techniques are also employed as the interaction prediction function in recommendation systems (He et al., 2017).
@@ 2.2. GCN-based Recommendation Models
Early recommendation approaches used random walks on graphs to leverage the high-order proximity between users and items, inferring indirect user preference (Christoffel et al., 2015; Fouss et al., 2007). Some recent approaches leverage the bipartite graph structure of user-item interactions to enhance these relationships (Yu et al., 2018). Furthermore, they probe further to uncover collaborative relations, such as similarities between users and items (Chen et al., 2019). The efficacy of these methods is closely tied to the quality of interactions sampled through the random walks, as they rely heavily on these samples for model training. In recent years, Graph Convolutional Networks (GCNs) have garnered significant attention in the realm of recommendation systems. Their rise in popularity can be attributed to their adeptness at learning node representations from non-Euclidean structures (van den Berg et al., 2018; Wang et al., 2019b, a, 2020; Huang et al., 2021; He et al., 2020; Liu et al., 2021; Wang and Leskovec, 2021; Lin et al., 2022; Cheng et al., 2023; Wei et al., 2023; Liu et al., 2023a). For example, NGCF (Wang et al., 2019b) propagates embeddings on the user-item interaction graph to harness the benefits of the high-order proximity. Drawing inspiration from studies focused on simplifying GCNs (Wu et al., 2019), researchers also critically scrutinize the intricacies of GCN-based recommendation models. Notably, He et al. (He et al., 2020) observed that the two commonly adopted design elements—feature transformation and nonlinear activation—did not contribute positively to the final performance. Consequently, they introduced LightGCN, a model that omits these two components and yields a marked improvement in recommendation performance. To achieve more efficient recommendations, the UltraGCN (Mao et al., 2021) model further simplifies graph convolutional networks by omitting embedding propagation layers. Additionally, it incorporates higher-order relationships through user-user graphs, leading to better performance and shorter running time compared to LightGCN. To further improve the neural graph collaborative filtering, NCL (Lin et al., 2022) explicitly incorporates the structural neighbors and semantic neighbors into contrastive pairs. Since the GCN-based recommendation methods suffer from the over-smoothing problem, IMP-GCN (Liu et al., 2021) refines user and item embeddings by conducting high-order graph convolutions within subgraphs. It can benefit from higher-order neighboring nodes by filtering out the noisy information from them. To learn more effective and robust representations, Fan et al. (Fan et al., 2022) proposed GTN, which can adaptively determine the reliability of interactions and effectively reduce unreliable interaction behaviors.
@@ 3.2. ClusterGCF Model
In this section, we present Cluster-based Graph Collaborative Filtering (ClusterGCF) model, along with its training methodology for recommendations. As illustrated in Fig. 2, ClusterGCF performs the first-order graph convolution over the original user-item interaction graph  [^2] Note that the first-order graph convolution is adopted as implemented in IMP-GCN (Liu et al., 2021). and the high-order graph convolution on the cluster-specific graphs to form the final representation of users and items. To construct cluster-specific graphs, we propose a soft node clustering approach to group users and items into multiple clusters. Next, based on the results of clustering, we assign a probability score to each node in the original user-item interaction graph for each cluster. This allows us to generate a graph that is specific to each cluster and shows the connections between users and items within that cluster. In the following, we will provide details of the ClusterGCF.
@@ 3.2.2. High-Order Graph Convolution
Cluster-Specific Graphs Combination. With the graph convolution operation, we can obtain the node representations for each layer of cluster-specific graphs. In other words, the messages are passed from the $k-$th layer neighboring nodes based on our constructed cluster-specific graphs. These messages are further combined as the $k-$th layer representations of user $u$ and item $i$ ($k\geq 2$):
@@
where $\bm{e_{u}^{(k)}}$ and $\bm{e_{i}^{(k)}}$ denote the features of user $u$ and item $i$ aggregated from the $k$-th layer neighbor nodes, respectively. Previous studies (He et al., 2020; Liu et al., 2021) have demonstrated the effectiveness of aggregating information from neighboring nodes through embedding summation. Thus, we apply the same method to aggregate the information within and between cluster-specific graphs. Note that we treat the information within different cluster-specific graphs equally because assigning different probabilities to the nodes in the different clusters can capture the varying importance.
@@
Due to its effectiveness, we adopt the same layer combination and prediction methodology as LightGCN (He et al., 2020). Specifically, after performing $k$ layers of graph convolution, the final embeddings of user $u$ and item $i$ are obtained by combing their embeddings across all layers of the ClusterGCF:
@@ 3.3. Model Training
In this work, our goal is to recommend a list of items to the target users by computing their preference score. For the optimization, we adopt the pairwise learning method, which is similar to the approach employed in previous works (Liu et al., 2021). To train our ClusterGCF model, we construct a triplet of $\{u,i^{+},i^{-}\}$, where $i^{+}$ represents a positive item and $i^{-}$ denotes a negative item. The objective function for ClusterGCF is defined as:
@@
Note that the user and item embeddings serve as inputs for both the node clustering and the recommendation system. We optimize these embeddings, which serve different purposes but are intricately linked within our optimization framework. Our approach ensures that the clustering process is continuously informed by the recommendation task, leading to clusters that are more relevant to the recommendation objective. Conversely, the embeddings benefit from the structure provided by the clusters, resulting in more robust and informative representations of users and items.